a) What is his marginal rate of substitution between X =3 and Y=4? If A=1, a = 0.4 and b = 0.5.


1
Expert's answer
2020-07-14T09:43:41-0400

Using given values, utility function is


U = AXaBy


U = X0.4Y0.5


"\text{MU}_{x}=\dfrac{\delta U}{\delta X}=\dfrac{0.4 \times 0.5Y}{0.6X}"


"\text{MU}_{y}=\dfrac{\delta U}{\delta X}=\dfrac{0.5 \times 0.4X}{0.5Y}"

"\text{Marginal rate of substitution} = \dfrac{- MUx }{ MUy}"


"=\dfrac{\dfrac{(0.4 x Y0.5) }{ X0.6}}{\dfrac{(0.5 x X0.4)}{ Y0.5}}"


"\dfrac{- (0.4)}{0.5} \times\dfrac{Y}{X}"



"\dfrac{ -4}{5} \times \dfrac{4}{3}"


"\dfrac{-16}{15}"


"\text{Marginal rate of substitution}= - 1.0667"